"You are told to build a rectangular fence along side a building. You have only 80 feet of fencing. The company wants the maximum amount of area enclosed. Find the dimensions that will lead to the most area"
All I need is the equation and I understand the problem from that point.

Answer by ankor@dixie-net.com(22740)

"You are told to build a rectangular fence along side a building.
You have only 80 feet of fencing.
The company wants the maximum amount of area enclosed.
Find the dimensions that will lead to the most area"
:
Along side a building, only 3 sides required, the equation:
L + 2W = 80
L = (80-2W)
:
Area = L * W
A = (80-2W)*W
A = -2W^2 + 80W
Find the axis of symmetry for max area